OpenStudy (anonymous):

sinx + cosx = (sinx/(1-cotx)) - (cosx/(tanx-1)) I'm supposed to balance the equations but i'm stuck... help please :

OpenStudy (mertsj):

\[\frac{\sin x}{1-\frac{\cos x}{\sin x}}-\frac{\cos x}{\frac{\sin x}{\cos x}-1}\]

OpenStudy (mertsj):

\[\frac{\sin ^2x}{\sin x-\cos x}-\frac{\cos ^2x}{\sin x-\cos x}\]

OpenStudy (mertsj):

\[\frac{\sin ^2x-\cos ^2x}{\sin x-\cos x}=\frac{(\sin x-\cos x)(\sin x+\cos x)}{\sin x-\cos x}\]

OpenStudy (mertsj):

\[\sin x+\cos x\]

OpenStudy (mertsj):

And so now the right side is the same as the left side and the identity is proved.

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

now which one do i give the best response lol